OpenStudy (anonymous):

Simplify. –7 + (–3)x + 5 + (–6)x A. –2 – 9x B. 2 – 3x C. 12 – 9x D. –12 – 3x

OpenStudy (anonymous):

I think it's A?

OpenStudy (anonymous):

Collect like terms, but yes you're correct.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

-7-3x+5-6x = -9x-2

OpenStudy (anonymous):

(-3x)+(-6x)=-9x -7+5=-2 so it is -2+(-9)x